a chemist has 300 grams of 20% hydrochloric acid solution. he wishes to drain some off and replace with an 80% solution so as to obtain a 25% solution. How many grams must he drain and replace with the 80% solution?
Answer by ankor@dixie-net.com(22740)

a chemist has 300 grams of 20% hydrochloric acid solution.
he wishes to drain some off and replace with an 80% solution so as to obtain a 25% solution.
How many grams must he drain and replace with the 80% solution?
:
Let x = the amt to be drained and the amt of 80% solution to be added
:
.20(300-x) + .80x = .25(300)
60 - .20x + .80x = 75
.60x = 75 - 60
.60x = 15
x = 15/.6
x = 25 grams drained, then 25 grams of 80% solution added
:
:
we can confirm this
.20(275) + .80(25) = .25(300)
55 + 20 = 75